# references.bib

### v1

@@ A.3.3 Seeker Reach-Avoid environment
Because of the dynamics of the seeker reach-avoid environment, we can simplify the relevant action set computation to the following optimization problem
$\!\max_{p}$ $\text{Vol}(\mathcal{A}^{r}_{k}(p))$ (37)
subject to $\mathcal{A}^{r}_{k}(p)\subseteq\mathcal{A}$
$\mathcal{S}^{r}_{k+\Delta t}=s_{k}\oplus\mathcal{A}^{r}_{k}(p)$
$\mathcal{S}^{r}_{k+\Delta t}\subseteq\left[-10,10\right]^{2}$
$\mathcal{S}^{r}_{k+\Delta t}\cap\mathcal{O}=\emptyset,$
Where $p$ is the parameter vector for the relevant set $\mathcal{A}^{r}(p)$, $\mathcal{S}_{1}$ is the reachable set of the agent in the next time step, $s_{0}\oplus\mathcal{A}^{r}(p)$ is the Minkowski sum (see 2.2) of the current state and the relevant action set, the set $\mathcal{O}$ represents the obstacle, and the box $\left[-10,10\right]^{2}$ is the outer boundary. We can solve this as a convex geometric program, using the same objective as in (36), and using the support function $\rho_{\mathcal{Z}}(l)=l^{T}c+\sum_{i=1}^{k}\big{|}l^{T}G_{(\cdot,i)}\big{|}$ of the zonotope $\langle c,G\rangle_{\mathcal{Z}}$ in the direction $l$:
$\!\max_{p}$ $\text{Vol}(\mathcal{A}^{r}_{k}(p))$ (38)
subject to $\|c^{\mathcal{A}^{r}_{k}}+\sum_{i}|G^{\mathcal{A}^{r}_{k}}_{(\cdot,i)}|\|_{\infty}\leq 1$
$\|c^{\mathcal{A}^{r}_{k}}+\sum_{i}-|G^{\mathcal{A}^{r}_{k}}_{(\cdot,i)}|\|_{\infty}\leq 1$
$\|c^{\mathcal{A}^{r}_{k}}+s_{k}+\sum_{i}|G^{\mathcal{A}^{r}_{k}}_{(\cdot,i)}|\|_{\infty}\leq 10$
$\|c^{\mathcal{A}^{r}_{k}}+s_{k}+\sum_{i}-|G^{\mathcal{A}^{r}_{k}}_{(\cdot,i)}|\|_{\infty}\leq 10$
$c^{\mathcal{A}^{r}_{k}}+s_{k}+\sum_{i}|n^{T}G^{\mathcal{A}^{r}_{k}}_{(\cdot,i)}|\leq b,$
The last constraint represents the containment of $S_{k+\Delta t}$ in the halfspace constructed by the normal vector $n$ at the intersection point of the line from the agent’s position to the center of the obstacle. More specifically, the halfspace is $\{x|n^{T}x\leq b\}$, where $n=\frac{o-s_{k}}{\|o-s_{k}\|}$, and $b=n^{T}(o-n)*r$.
@@ A.4 Computation of zonotope boundary points
The boundary point $p\in\mathbb{R}^{N}$ on a zonotope $\langle c,G\rangle_{\mathcal{Z}}\subset\mathbb{R}^{N}$ in a direction $d\in\mathbb{R}^{N}$ starting from an origin point $x\in\mathbb{R}^{N}$ is obtained by solving the linear program
@@
A.5 Parameters for environments
[…]
@@ A.5 Parameters for environments
Template matrix $\tilde{G}$ [[1, 1, 1, 0], [1, -1, 0, 1]] [[1, 1, 1], [1, -1, 0]] $I_{4}$
A.6 Computational runtime for training
Note that the increase of runtime between the masking approaches and baseline is mainly due to the computation of the relevant action sets. Since the relevant action set is a zonotope, the generator mask does not require expensive additional computations. The increased runtime for the ray mask mainly originate from the computation of the boundary point (see Appendix A.4). For the distributional mask the increased runtime is mostly caused by the Markov-Chain Monte-Carlo sampling of the action.
Table 3: Mean and standard deviation of runtime in hours for training runs on utilized machine.
Baseline Ray mask Generator mask Distributional mask
Seeker $0.045\pm 0.001$ $0.865\pm 0.013$ $0.699\pm 0.016$ $1.62\pm 0.017$
2D Quadrotor $1.049\pm 0.197$ $3.140\pm 0.097$ $2.533\pm 0.021$ $4.083\pm 0.080$
3D Quadrotor $0.493\pm 0.077$ $2.031\pm 0.258$ $1.528\pm 0.010$ $3.130\pm 0.045$
A.7 Hyperparameters for learning algorithms
We specify the hyperparameters for the three approaches and baseline in the Seeker environment (Table 4), the 2D Quadrotor (Table 5, and the 3D Quadrotor environment (Table 6). These were achieved through hyperparameter optimization with 50 trials using the Tree-structured Parzen estimator of optuna [akiba2019optuna].
@@ A.7 Hyperparameters for learning algorithms
Parameter Baseline Ray mask Generator mask Distributional mask
Learning rate $55.432\mathrm{E}-$ 5 $48.249\mathrm{E}-$ 4 $43.446\mathrm{E}-$ 4 $53.848\mathrm{E}-$ 5
Discount factor $\gamma$ $0.98$ $0.98$ $0.98$ $0.98$
Steps per update $32$ $256$ $2084$ $32$
Optimization epochs $4$ $8$ $16$ $4$
Minibatch size $8$ $128$ $256$ $8$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $54.709\mathrm{E}-$ 5 $71.657\mathrm{E}-$ 7 $76.609\mathrm{E}-$ 7 $63.326\mathrm{E}-$ 6
Initial log stddev. $-1.183$ $-0.010$ $-0.255$ $-1.213$
Value function coefficient $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.9$ $0.9$ $0.9$ $0.9$
@@
Neurons per layer $32$ $32$ $32$ $32$
[…]
Parameter Baseline Ray mask Generator mask Distributional mask
Learning rate $41.240\mathrm{E}-$ 4 $47.919\mathrm{E}-$ 4 $34.342\mathrm{E}-$ 3 $43.94\mathrm{E}-$ 4
Discount factor $\gamma$ $0.99$ $0.99$ $0.99$ $0.99$
Steps per update $256$ $1024$ $1024$ $1024$
Optimization epochs $32$ $8$ $8$ $8$
Minibatch size $64$ $128$ $128$ $128$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $28.9\mathrm{E}-$ 2 $25.645\mathrm{E}-$ 2 $25.077\mathrm{E}-$ 2 $35.99\mathrm{E}-$ 3
Initial log stddev. $-0.4366$ $-0.7838$ $-1.2509$ $-1.217$
Value function coefficient $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.95$ $0.95$ $0.95$ $0.95$
Activation function ReLU ReLU ReLU ReLU
Hidden layers $2$ $2$ $2$ $2$
Neurons per layer $256$ $256$ $256$ $256$
[…]
Parameter Baseline Ray mask Generator mask Distributional mask
Learning rate $42.384\mathrm{E}-$ 4 $31.077\mathrm{E}-$ 3 $59.241\mathrm{E}-$ 5 $47.875\mathrm{E}-$ 4
Discount factor $\gamma$ $0.98$ $0.98$ $0.98$ $0.98$
Steps per update $32$ $128$ $128$ $64$
Optimization epochs $8$ $4$ $16$ $4$
Minibatch size $16$ $32$ $16$ $64$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $55.847\mathrm{E}-$ 5 $71.136\mathrm{E}-$ 7 $73.408\mathrm{E}-$ 7 $62.746\mathrm{E}-$ 6
Initial log stddev. $-3.609$ $-1.793$ $-1.363$ $-1.880$
Value function coefficient $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.9$ $0.9$ $0.9$ $0.9$
Activation function ReLU ReLU ReLU ReLU
Hidden layers $2$ $2$ $2$ $2$
Neurons per layer $32$ $32$ $32$ $32$

### v2

@@ A.3.3 Seeker Reach-Avoid environment
Because of the dynamics of the Seeker Reach-Avoid environment, we can simplify the relevant action set computation to the following optimization problem:
$\!\max_{p}$ $\widetilde{\texttt{Vol}}(\mathcal{A}^{r}(p))$ (35)
subject to $\mathcal{A}^{r}(p)\subseteq\mathcal{A}$
$\mathcal{S}^{r}_{\Delta t}=s\oplus\mathcal{A}^{r}(p)$
$\mathcal{S}^{r}_{\Delta t}\subseteq\left[-10,10\right]^{2}$
$\mathcal{S}^{r}_{\Delta t}\cap\mathcal{O}=\emptyset,$
where $\mathcal{S}_{\Delta t}$ is the reachable set of the agent in the next time step, the set $\mathcal{O}$ represents the obstacle, and the box $\left[-10,10\right]^{2}$ is the outer boundary of the state space. The constraints represent three intuitive geometric constraints. First, containment of the relevant action set in the action set (usually, the interval $[-1,1]^{2}$, second, the containment of the relevant state set of the next time step in the environment boundaries, and the last constraint enforces that there is no collision with the obstacle possible. To enforce these constraints for zonotopes, we use the support function [Althoff2016supportfunctions, Lemma 1]
$\rho_{\mathcal{Z}}(l)=l^{T}c+\sum_{i=1}^{P}\big{|}l^{T}G_{(\cdot,i)}\big{|},$ (36)
where $l$ is the vector in the direction of interest. For the first constraint, we need to ensure that the support function for the basis vectors $e_{1}$, and $e_{2}$ (and there negative versions) are less than or equal to $1$, i.e.,
$\begin{bmatrix}\rho_{\mathcal{A}^{r}}(e_{1})\\ \rho_{\mathcal{A}^{r}}(e_{2})\end{bmatrix}=\begin{bmatrix}c_{1}^{\mathcal{A}^{r}}+\sum_{i=1}^{P}\big{|}G_{(1,i)}^{\mathcal{A}^{r}}\big{|}\\ c_{2}^{\mathcal{A}^{r}}+\sum_{i=1}^{P}\big{|}G_{(2,i)}^{\mathcal{A}^{r}}\big{|}\end{bmatrix}=c^{\mathcal{A}^{r}}+\sum_{i=1}^{P}\big{|}G_{(\cdot,i)}^{\mathcal{A}^{r}}\big{|}\leq\begin{bmatrix}1\\ 1\end{bmatrix}.$ (37)
From (36) it is apparent that using the negative basis vectors $-e_{1}$ and $-e_{2}$, just flips the sign of $c^{\mathcal{A}^{r}}$ in (37). The constraints for $\mathcal{S}^{r}_{\Delta t}\subseteq\left[-10,10\right]^{2}$ can be constructed similarly. For simplicity, we express the element-wise inequality in (37) using the $L_{\infty}$-norm, and write the optimization problem as
$\!\max_{p}$ $\widetilde{\texttt{Vol}}(\mathcal{A}^{r}(p))$ (38)
subject to $\|c^{\mathcal{A}^{r}}+\sum_{i}|G^{\mathcal{A}^{r}}_{(\cdot,i)}|\|_{\infty}\leq 1$
$\|-c^{\mathcal{A}^{r}}+\sum_{i}|G^{\mathcal{A}^{r}}_{(\cdot,i)}|\|_{\infty}\leq 1$
$\|c^{\mathcal{A}^{r}}+s+\sum_{i}|G^{\mathcal{A}^{r}}_{(\cdot,i)}|\|_{\infty}\leq 10$
$\|-c^{\mathcal{A}^{r}}+s+\sum_{i}|G^{\mathcal{A}^{r}}_{(\cdot,i)}|\|_{\infty}\leq 10$
$n^{T}(c^{\mathcal{A}^{r}}+s)+\sum_{i}|n^{T}G^{\mathcal{A}^{r}}_{(\cdot,i)}|\leq b.$
The last constraint represents an under-approximation of $\mathcal{S}^{r}_{\Delta t}\cap\mathcal{O}=\emptyset$ enforcing the containment of $\mathcal{A}^{r}$ in the halfspace $\{x|n^{T}x\leq b\}$ through the support function $\rho_{\mathcal{A}^{r}}(n)$. The directional vector $n=\frac{o-s}{\|o-s\|}$ where $o$ is the center of the obstacle, and the offset $b=n^{T}(o-n)r$ where $r$ is the radius of the obstacle. This ensures that the intersection between the reachable set $\mathcal{S}^{r}_{\Delta t}$ and the obstacle $\mathcal{O}$ will be empty since the halfspace is constructed tangential to the obstacle at the intersection point of the line from the agent to the center of the obstacle.
@@ A.4 Computation of zonotope boundary points
The boundary point $p\in\mathbb{R}^{N}$ on a zonotope $\langle c,G\rangle_{\mathcal{Z}}\subset\mathbb{R}^{N}$ in a direction $d\in\mathbb{R}^{N}$ starting from a point $x\in\mathbb{R}^{N}$ is obtained by solving the linear program
@@
A.5 Parameters and dynamics for environments
We provide the action space bounds, the state space bounds, and the generator template matrix $\tilde{G}$ for the Seeker and Quadrotor environments in Table 2.
@@ A.5 Parameters and dynamics for environments
Template matrix $\tilde{G}$ $\begin{bmatrix}1&1&1&0\\ 1&-1&0&1\end{bmatrix}$ $\begin{bmatrix}1&1&1\\ 1&-1&0\end{bmatrix}$ $I_{4}$
The system dynamics of the 2D Quadrotor is:
@@
The 3D Quadrotor is modeled in a twelve-dimensional state space with state $s=[x,y,z,\dot{x},\dot{y},\dot{z},\theta,\phi,\psi,\dot{\theta},\dot{\phi},\dot{\psi}]^{T}$ and four-dimensional action space with action $a=[a_{1},a_{2},a_{3},a_{4}]^{T}$. The system dynamics are $\dot{s}=[\dot{x},\dot{y},\dot{z},-9.81\phi,9.81\theta,a_{1},\dot{\theta},\dot{\phi},\dot{\psi},a_{2},a_{3},a_{4}]^{T}$ [Kaynama2015].
For the Walker2D environment, we use the standard parameters of the gymnasium implementation [^1] https://gymnasium.farama.org/environments/mujoco/walker2d/. The relevant action set is a zonotope under-approximation of the relevant action set stated in (20) with 36 generators and $\alpha_{P}=1$.
A.6 Computational time for training
Note that the increased computation time of the masking approaches compared to the baseline is mainly due to the computation of the relevant action sets. Since the relevant action set is a zonotope, the generator mask does not require expensive additional computations. The increased runtime for the ray mask mainly originates from the computation of the boundary points (see Appendix A.4). For the distributional mask, the increased runtime is mostly caused by the Markov chain Monte Carlo sampling of the action.
Table 3: Mean and standard deviation of runtime in hours for training runs on the utilized machine.
Seeker 2D Quadrotor 3D Quadrotor Walker2D
Baseline $0.045\pm 0.001$ $1.049\pm 0.197$ $0.493\pm 0.077$ $0.350\pm 0.007$
Ray mask $0.865\pm 0.013$ $3.140\pm 0.097$ $2.031\pm 0.258$ $0.962\pm 0.043$
Generator mask $0.699\pm 0.016$ $2.533\pm 0.021$ $1.528\pm 0.010$ $0.557\pm 0.021$
Distributional mask $1.620\pm 0.017$ $4.083\pm 0.080$ $3.130\pm 0.045$ $\approx 60$
Replacement $0.823\pm 0.015$ $2.782\pm 0.048$ $1.765\pm 0.033$ $0.880\pm 0.039$
A.7 Quantitative results for Walker2D environment compared to standard PPO
If we remove the power constraint that resets the environment whenever an action outside of $\mathcal{A}^{r}$ is selected, the standard PPO is able to learn a policy that performs slightly better than the ray mask towards the end of training while converging slightly slower (see Figure 4). However, during deployment, only $0.05\%$ of the actions from this policy are from within $\mathcal{A}^{r}$. Yet, the reward that standard PPO achieves during deployment is on average $2177.57$ with standard deviation $1238.42$. This is slightly higher than the deployment results for the ray mask (see Table 1).
Figure 4: Average reward curves for Walker2D with transparent bootstrapped 95% confidence interval including standard PPO as additional baseline.
A.8 Qualitative results for the Seeker environment
Figure 5 presents ten example trajectories for a trained agent using each of the five approaches under consideration. Notably, the continuous action masking agents solve the task by safely reaching the goal while the PPO baseline still collides with the obstacle. The replacement approach is ineffective, with the agent frequently failing to reach the goal. In fact, for the ten displayed rollouts, no replacement agent reaches the goal. This discrepancy with respect to the training performance where the replacement agent reaches the goal frequently is due to using the policy in a deterministic setting for deployment. The simple dynamics of the Seeker environment make it possible to directly visualize the relevant action set $\mathcal{A}^{r}$ at each time step along the trajectory, which is depicted in the lower half of Figure 5. For the generator mask, the least amount of $\mathcal{A}^{r}$ is plotted, which indicates that the generator mask agent reaches the goal most efficiently, i.e., the lowest amount of steps are required.
Figure 5: Qualitative deployment results for ten initial states and one goal-obstacle configuration for the Seeker environment. The top half shows ten trajectories with randomly sampled starting states. The lower half depicts the relevant action set (green polygon) for each time step along one trajectory.
A.9 Hyperparameters for learning algorithms
We specify the hyperparameters for the three masking approaches, replacement and baseline in the Seeker (Table 4), the 2D Quadrotor (Table 5), and the 3D Quadrotor environment (Table 6). These were obtained through hyperparameter optimization with 50 trials using the tree-structured Parzen estimator of Optuna with the default parameters [akiba2019optuna].
[…]
Parameter Baseline Ray Generator Distributional Replacement
Learning rate $55.43\mathrm{E}-$ 5 $48.25\mathrm{E}-$ 4 $43.45\mathrm{E}-$ 4 $53.85\mathrm{E}-$ 5 $61.92\mathrm{E}-$ 6
Discount factor $\gamma$ $0.98$ $0.98$ $0.98$ $0.98$ $0.98$
Steps per update $32$ $256$ $2084$ $32$ $128$
Optimization epochs $4$ $8$ $16$ $4$ $4$
Minibatch size $8$ $128$ $256$ $8$ $128$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $54.71\mathrm{E}-$ 5 $71.66\mathrm{E}-$ 7 $76.61\mathrm{E}-$ 7 $63.33\mathrm{E}-$ 6 $71.83\mathrm{E}-$ 7
Initial log std dev $-1.183$ $-0.010$ $-0.255$ $-1.213$ $-1.064$
Value function coeff. $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.9$ $0.9$ $0.9$ $0.9$ $0.9$
Activation function ReLU ReLU ReLU ReLU ReLU
Hidden layers $2$ $2$ $2$ $2$ $2$
Neurons per layer $32$ $32$ $32$ $32$ $32$
[…]
Parameter Baseline Ray Generator Distributional Replacement
Learning rate $41.24\mathrm{E}-$ 4 $47.92\mathrm{E}-$ 4 $34.34\mathrm{E}-$ 3 $43.94\mathrm{E}-$ 4 $41.13\mathrm{E}-$ 4
Discount factor $\gamma$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Steps per update $256$ $1024$ $1024$ $1024$ $512$
Optimization epochs $32$ $8$ $8$ $8$ $8$
Minibatch size $64$ $128$ $128$ $128$ $128$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $28.9\mathrm{E}-$ 2 $25.65\mathrm{E}-$ 2 $25.08\mathrm{E}-$ 2 $35.99\mathrm{E}-$ 3 $35.42\mathrm{E}-$ 3
Initial log std dev $-0.437$ $-0.784$ $-1.251$ $-1.217$ $-1.019$
Value function coeff. $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.95$ $0.95$ $0.95$ $0.95$ $0.95$
Activation function ReLU ReLU ReLU ReLU ReLU
Hidden layers $2$ $2$ $2$ $2$ $2$
Neurons per layer $256$ $256$ $256$ $256$ $256$
[…]
Parameter Baseline Ray Generator Distributional Replacement
Learning rate $42.38\mathrm{E}-$ 4 $31.08\mathrm{E}-$ 3 $59.24\mathrm{E}-$ 5 $47.88\mathrm{E}-$ 4 $46.25\mathrm{E}-$ 4
Discount factor $\gamma$ $0.98$ $0.98$ $0.98$ $0.98$ $0.98$
Steps per update $32$ $128$ $128$ $64$ $128$
Optimization epochs $8$ $4$ $16$ $4$ $4$
Minibatch size $16$ $32$ $16$ $64$ $64$
Max gradient norm $0.9$ $0.9$ $0.9$ $0.9$ $0.9$
Entropy coefficient $55.85\mathrm{E}-$ 5 $71.14\mathrm{E}-$ 7 $73.41\mathrm{E}-$ 7 $62.75\mathrm{E}-$ 6 $61.88\mathrm{E}-$ 6
Initial log std dev $-3.609$ $-1.793$ $-1.363$ $-1.880$ $-1.582$
Value function coeff. $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Clipping range $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
GAE $\lambda$ $0.9$ $0.9$ $0.9$ $0.9$ $0.9$
Activation function ReLU ReLU ReLU ReLU ReLU
Hidden layers $2$ $2$ $2$ $2$ $2$
Neurons per layer $32$ $32$ $32$ $32$ $32$
[…]
Parameter Baseline Replacement Ray mask Generator mask
Learning rate $56.992\mathrm{E}-$ 5 $34.700\mathrm{E}-$ 3 $42.607\mathrm{E}-$ 4 $41.719\mathrm{E}-$ 4
Discount factor $\gamma$ $0.99$ $0.99$ $0.99$ $0.99$
Steps per update $2048$ $2048$ $2048$ $2048$
Minibatch size $128$ $64$ $16$ $32$
Max gradient norm $0.603$ $0.336$ $0.156$ $0.152$
Entropy coefficient $76.559\mathrm{E}-$ 7 $65.960\mathrm{E}-$ 6 $64.992\mathrm{E}-$ 6 $57.488\mathrm{E}-$ 5
Clipping range $0.165$ $0.131$ $0.102$ $0.192$
GAE $\lambda$ $0.970$ $0.944$ $0.919$ $0.957$
Value function coefficient $0.259$ $0.330$ $0.181$ $0.500$
@@ A.9 Hyperparameters for learning algorithms
Neurons per layer $64$ $64$ $64$ $64$